I could use some help on this problem.
Which of these is the first step in solving x^2 + 6x + 1 = 0 by completing the square?
a. (x + 6)^2 = -1
b. x^2 + 6x = -1
c. x^2 + 6x + 9 - 9 + 1 = 0
d. (x + 6)^2 + 1 = 1
I came up with A., but I'm not sure if that is correct.

Found 3 solutions by Alan3354, scott8148, stanbon:
Answer by Alan3354(69443)

Which of these is the first step in solving x^2 + 6x + 1 = 0 by completing the square?
a. (x + 6)^2 = -1
b. x^2 + 6x = -1
c. x^2 + 6x + 9 - 9 + 1 = 0
d. (x + 6)^2 + 1 = 1
I came up with A., but I'm not sure if that is correct.
------------
B is the next thing to do, taking it one step at a time.

Answer by scott8148(6628)

it's b

the next steps are:

x^2 + 6x + (6/2)^2 = -1 + (6/2)^2

(x + 3)^2 = 8

x + 3 = ± 2√2

x = -3 ± 2√2

Answer by stanbon(75887)

Which of these is the first step in solving x^2 + 6x + 1 = 0 by completing the square?
---------
x^2 + 6x + 1 = 0
----
x^2+6x+(3^2) - (3^2) + 1 = 0
------
(x+3)^2 - 9 + 1 = 0
----
(x+1)^2 = 8
etc.
=====================================
Cheers,
Stan H.
==================


a. (x + 6)^2 = -1
b. x^2 + 6x = -1
c. x^2 + 6x + 9 - 9 + 1 = 0
d. (x + 6)^2 + 1 = 1